OpenStudy (jojo12g):

please help me im stuck and i need your help A spinner has 9 equal sections labeled 1 through 9. What is the probability of spinning an odd number on the spinner and tossing heads on a coin? A.11/15 B.1/4 C.1/6 D.2/15

OpenStudy (jojo12g):

i will fan and medal please heelp noone else has

OpenStudy (jojo12g):

i think d

OpenStudy (nikato):

none of the answers seem right to me

OpenStudy (jojo12g):

can u help with anoter please

OpenStudy (nikato):

because prob. of spinning an odd= 5/9 and having tail si 1/2 so (5/9)x(1/2)= 5/18

OpenStudy (nikato):

and sure

OpenStudy (jojo12g):

A bag contains 2 gold coins and 5 silver coins. A number cube has the numbers 1 through 6 on its faces. What is the probability of drawing a silver coin from the bag and rolling an even number on the number cube? A.5/42 B. 5/14 C. 5/9 D.2/3

OpenStudy (nikato):

so whats the probability of picking a silver coin?

OpenStudy (jojo12g):

1/6

OpenStudy (nikato):

not really. a) how many silver coins are there? b) how many coins total are in the bag?

OpenStudy (jojo12g):

5 silver and 6

OpenStudy (nikato):

yes 5 silver but the total is actually 7. 5 silve and 2 gold so 5+2=7

OpenStudy (nikato):

so the proability of picking a silver coin is 5/7 ok?

OpenStudy (jojo12g):

k

OpenStudy (jojo12g):

simplify

OpenStudy (jojo12g):

nevermind you didnt even help

OpenStudy (nikato):

wait, now we have to find the probability of rolling an even number

OpenStudy (nikato):

there are 3 even number in total. and 6 possible out comes so 3/6

OpenStudy (nikato):

now multiply them together